OpenStudy (anonymous):

Kramers rule 2x+7y=-4 X+4y=-3 How is it done?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Kramer's Rule 2x+7Y=-4 X+4+-3 how is it done

OpenStudy (radar):

You can do the simultaneous equation. Do this multiply the second equation by 2 getting: 2x+8y=-6, then subtract the first equation 2x+7y=-4 y = -2 x would be 5

OpenStudy (radar):

Sorry you wanted to use matrice, that was not Cramer's Rule lol. But for only 2 or 3 linear equations, it would suffice. suggest you google cramers rule.